The sum of two consecutive odd integers is 60. State the smaller odd integer x
Found 2 solutions by Alan3354, josgarithmetic:
Answer by Alan3354(69443)

The sum of two consecutive odd integers is 60. State the smaller odd integer x
-----------
60/2 = 30
--> 29 & 31
================
Or,
---
n-1 + n+1 = 60
2n = 60
n = 30
etc
=================
Or,
n + n+2 = 60
etc

Answer by josgarithmetic(39799)

29

( No algebra was used.)